a rectangle sand box is 1.75 feet longer than it is wide. if the perimeter is 37.5ft what are the dimensions of the sandbox?
Answer by CubeyThePenguin(3113)

width = w
length = w + 1.75

2(w) + 2(w + 1.75) = 37.5
4w + 3.5 = 37.5
4w = 34
w = 34/4 = 8.5

dimensions: 8.5 ft by 10.25 ft